QUESTION # 1

The method of tree-ring dating gave the following years A.D. for an archaeological excavation site. Assume that the population of x values has an approximately normal distribution.

1,264 1,278 1,208 1,250 1,268 1,316 1,275 1,317 1,275

(a)

Use a calculator with mean and standard deviation keys to find the sample mean year x and sample standard deviation s. (Round your answers to the nearest whole number.)

x= ________-A.D.s= yr

(b)Find a 90% confidence interval for the mean of all tree-ring dates from this archaeological site. Round your answers to the nearest whole number.

lower limit__________ A.D

.upper limit__________ A.D.

QUESTION # 2

What percentage of hospitals provide at least some charity care? Based on a random sample of hospital reports from eastern states, the following information is obtained (units in percentage of hospitals providing at least some charity care):

56.9 55.5 53.2 65.5 59.0 64.7 70.1 64.7 53.5 78.2

Assume that the population of x values has an approximately normal distribution.

(a) Use a calculator with mean and sample standard deviation keys to find the sample mean percentage x and the sample standard deviation s. (Round your answers to one decimal place.)

x = _______%
s = _______%


(b) Find a 90% confidence interval for the population average μ of the percentage of hospitals providing at least some charity care. (Round your answers to one decimal place.)

lower limit     _______ %
upper limit     _______%
Answer #1

1. a)  LT = 1272 A.D

s=33 yr

b) Critical value te = t8.0.1/2 = 1.860

Margin of error, E = t_{c}\frac{s}{\sqrt{n}} =  1.86053 = 20.46

Lower Limit = T - E= 1272 - 20.46 = 1251 A.D

Upper Limit = T + E = 1272 + 20.46 = 1292 A. D

2. a)  I= 62.1 %

s = 8,0 %

b) Critical value te = t9,0.1/2 = 1.833

Margin of error, E = t_{c}\frac{s}{\sqrt{n}} =  1.833- = 4.6

Lower Limit = \bar{x} - E= 62.1 -4.6 = 57.5 %

Upper Limit = T - E= 62.1 + 4.6 = 66.7 %